8. A projectile ia fired aa seen below. Negleot air reaiatance, use the given information and g- 9.80 m/a2 a Caloulate the initial apeed of the projeotile. Anawer: 12.0 m/s Suggeation: Start with the trajeotory equation and solve for initial apeed. b. Caloulate the height the projeotile reachea above the s-axia. o. Caloulate the projeotilea total time of flight. Anawer: 3.12 a y-axis 60 x-axis 15.24 m - 18.70 m.
Answer #1

a)

Considering motion along horizontal

R = vo cos x * t

18.7 = vo* cos 60 * t

vo* t = 37.4 ... (I)

Considering motion along vertical

0 = 15.24 + vo* sin 60* t - 0. 5* * 9.8* t^2

0 = 15.24 + 37.4* sin 60 - 4.9 t^2

t = 3.118 s

Putting in (i)

Vo = 12 m/s

=========

b)

Height above x axis

H = (Vo sin 60) ^2 / ( 2g)

H = (12 * sin 60)^2 / 19.6

H = 5.51 m

=======

C)

T = 3.118 ≈ 3.12 seconds

========

Do comment in case any doubt, will reply for sure. Goodluck